OpenStudy (anonymous):

What is 3(x-2)^2+7=154 ?

OpenStudy (anonymous):

firstly I'd expand the brackets, so you'd get 3(x^2+4-4x) + 7=154 then multiply the brackets, 3x^2 + 12 -12x + 7 =154 simplify => 3x^2 -12x -135 = 0 there are no factor for that quadratic equation so we use the quadratic formula therefore: x = {-(-12) (+/-) the squareroot of [ (-12)^2 - 4(3)(-135)]} all divided by (2*3) by punching into a calculator, you'd get one solution which is 9 but remember to take into the negative sign of the squareroot as well, so the other solution would be -5. overall, x= -5 and x=9